OpenStudy (anonymous):

okay module two quizz in algebra 1. please only question 4

OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (michele_laino):

please, you have to solve two equation: case 1) if x-0.525>=0 then you have to solve this: \[x-0.525=0.01\] whereas case2) if x-0.525<0, yiou have to solve this: \[-(x-0.525)=0.01\]

OpenStudy (michele_laino):

case 1) x=0.525+0.01=0.535, case 2) \[-x+0.525=0.01\] \[x=0.525-0.01=0.515\]

OpenStudy (anonymous):

thanks for ur help(: i got it

OpenStudy (michele_laino):

thank you!